Find a size change that maps f(x)=ax^2 onto g(x)=bx^2 where a and b are non-zero real constants.
Answer by Alan3354(69443)

Find a size change that maps f(x)=ax^2 onto g(x)=bx^2 where a and b are non-zero real constants.
--------
b/a
---
ax^2*(b/a) = bx^2